# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Uncertain Dynamics

### v2

@@ Abstract
Index Terms:
Collision avoidance, integrated planning and control, planning under uncertainty, reinforcement learning.
@@ III-A Composite Vector Field
where $\hat{(\cdot)}$ is the normalization notation (i.e., for a nonzero vector $v\in\mathbb{R}^{n}$, $\hat{v}=v/\lVert v\rVert$), $\sqcup_{\mathcal{Q}}(\xi)=\frac{f_{1}(\xi)}{f_{1}(\xi)+f_{2}(\xi)},\sqcap_{\mathcal{R}}(\xi)=\frac{f_{2}(\xi)}{f_{1}(\xi)+f_{2}(\xi)}$ are smooth bump functions, where $f_{1}(\xi)=0$ if $\varphi(\xi)\leq c$ and $f_{1}(\xi)=\exp\left(\frac{l_{1}}{c-\varphi(\xi)}\right)$ if $\varphi(\xi)>c$, $f_{2}({\xi})=\exp\left(\frac{l_{2}}{\varphi(\xi)}\right)$ if $\varphi(\xi)<0$ and $f_{2}({\xi})=0$ if $\varphi(\xi)\geq 0$, and $l_{1},l_{2}>0$ are coefficients for changing the decaying or increasing rate. Note that for simplicity, the subscripts $i$ of related symbols have been omitted above. These smooth bump functions blend parts of different vector fields and create a composite vector field for path following and collision avoidance; for more details, see [5]. To understand the composite vector field intuitively, we illustrate the composite vector field (1) in Fig. 3. For $\chi(\xi)$ in (1), it is equal to $\hat{\chi}_{\mathcal{P}}(\xi)$, $\sqcup_{\mathcal{Q}}\left(\xi\right)\hat{\chi}_{\mathcal{P}}\left(\xi\right)+\sqcap_{\mathcal{R}}\left(\xi\right)\hat{\chi}_{\mathcal{R}}\left(\xi\right)$, and $\hat{\chi}_{\mathcal{R}}(\xi)$ within the three regions ${}^{\rm{ex}}{\mathcal{R}}$, ${}^{\rm{ex}}{\mathcal{Q}}\cap{}^{\mathrm{in}}\mathcal{R}$, and ${}^{\mathrm{in}}\mathcal{Q}$, respectively.
@@
$\mathcal{C}_{\mathcal{P}}$ =$=\left\{\xi\in\mathbb{R}^{2}:\chi_{\mathcal{P}}(\xi)=0\right\}=\left\{\xi\in\mathbb{R}^{2}:\nabla\phi(\xi)=0\right\},$
$\mathcal{C}_{\mathcal{R}_{i}}$ =$=\left\{\xi\in\mathbb{R}^{2}:\chi_{\mathcal{R}_{i}}(\xi)=0\right\}=\left\{\xi\in\mathbb{R}^{2}:\nabla\varphi_{i}(\xi)=0\right\}.$
@@ III-B Offline Deep Koopman Operator-based System Modeling
$\varUpsilon_{\infty}(x_{k+1})=(\mathcal{K}\varUpsilon_{\infty})(x_{k},u_{k}),$ (3)
@@ III-D2 Optimal Motion Planning to Avoid Moving Obstacles
$\varUpsilon(x_{k+1})=\underbrace{A\varUpsilon(x_{k})+Bu_{k}}_{f_{\text{nom}}(\varUpsilon(x_{k}),u_{k})}+B_{s}\underbrace{\left(g(\varUpsilon(x_{k}),u_{k})+w_{k}\right)}_{y_{k}},$ (9)
@@
$\tilde{x}_{k+1}=A\tilde{x}_{k}+B\tilde{u}_{k}+B_{s}\underbrace{\left(g(\varUpsilon(x_{k}),u_{k})+w_{k}\right)}_{y_{k}},$ (11)
where $\tilde{x}_{k}=\varUpsilon(x_{k})-\varUpsilon(x_{r,k})$ is the error state, $x_{r,k}$ is the reference state, and $\tilde{u}_{k}=u_{k}-u_{r,k}$ is the control input.
@@
$V_{\infty}(\tilde{x}_{k})=\sum_{\tau=k}^{\infty}\gamma^{\tau-k}L(\tilde{x}_{\tau},\tilde{u}_{\tau}),$ (12)
where $0<\gamma\leq 1$, $L(\tilde{x}_{\tau},\tilde{u}_{\tau})=\tilde{x}_{\tau}^{\top}Q\tilde{x}_{\tau}+\tilde{u}_{\tau}^{\top}R\tilde{u}_{\tau}$ is the cost function, $Q\succeq 0\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$ is positive semi-definite, and $R\succ 0\in\mathbb{R}^{n_{u}\times n_{u}}$ is positive definite. The second subproblem is how the robot can avoid moving obstacles. Combining the two subproblems, we formulate the optimal OMP as below:
@@ IV-A1 Composite Vector Field with Kinodynamic Constraints
[…]
@@
where the adjustable coefficient $k_{c,i}>0$ is used to change the convergence rate to $\mathcal{Q}_{\rm{vrtl}}$, and $\exp(\cdot)$ denotes the exponential function. Thus, in this design, $10<{s_{i}}(\xi)<$ 1 if $\xi\in\mathcal{M}_{b}$; ${s_{i}}(\xi)=1$, otherwise.
@@ IV-A3 Discrete-time Kinodynamic Guiding Vector Field
where $\xi_{0}\in\mathbb{R}^{2}$ is the initial robot position and $\beta$ is the step length, and $\xi_{k,grid}=\arg\min_{\xi_{g}\in\mathcal{G}}||\xi_{k}-\xi_{g}||$, where the set $\mathcal{G}\subseteq\mathbb{R}^{2}$ is the grid map consisting of a finite number of chosen points in a selected region. In terms of computational efficiency, it is advisable to precompute $\chi_{c}(\xi)$ on a mesh grid map. This makes it possible to quickly identify the nearest vector $\chi_{c}(\xi_{k,grid})$ on the mesh grip map based on the current position $\xi_{k}$. Finally, we can obtain the planned trajectory. Note that a vector with nearly zero norms is not selected as the current vector but rather inherits the vector from the previous state to avoid suffering from the singularity issue. Namely, if $||\chi_{c}(\xi_{k,grid})||<\epsilon$, then $\chi_{c}(\xi_{k,grid})\leftarrow\chi_{c}(\xi_{k-1,grid})$, where $\epsilon$ is a small positive number. This allows us to plan guiding trajectories directly without causing the deadlock problem from singular points.
@@ IV-B Online Compensation to Update the Offline Trained Deep Koopman Model
${d}(z_{k})=B_{d}^{\dagger}(\varUpsilon({x}_{k+1})-f_{\text{nom}}\left(\varUpsilon({x}_{k}),{u}_{k}\right)),$ (18)
@@
$\varUpsilon({x}_{k+1})=f_{\text{nom}}(\varUpsilon({x}_{k}),{u}_{k})+{d}(z_{k})+\epsilon,$ (19)
where $\epsilon\in\mathbb{R}^{n_{\mathcal{K}}}$ is the estimation error. The reference trajectory can be expressed as $\varUpsilon(x_{r,k+1})=f_{\text{nom}}(\varUpsilon({x}_{r,k}),{u}_{r,k})$.
We can obtain the Jacobian matrix of (19) at a reference state $(\varUpsilon({x}_{r,k}),{u}_{r,k})$, i.e.,
@@
$\hat{\tilde{x}}_{k+1}=f_{\text{nom}}(\tilde{x}_{k},\tilde{u}_{k})+d({z_{k}}).$ (22)
@@ IV-C Exponential Barrier Function Incorporated Cost Function to Avoid Moving Obstacles
Due to the adoption of rule-based switching between tracking control and obstacle-avoiding modes in [9], without considering obstacle velocities, it does not guarantee the safety of mobile robots under extreme conditions such as the hazardous driving behavior of opposing vehicles. Our idea stems from the work[37] in completing suicidal game tasks, where the “evader” aspires to avoid being caught by the “pursuer” by computing safe actions. Similarly, when moving obstacles obstruct the road ahead, our strategy is to keep a safe distance from obstacles and track the desired trajectory if the ego robot cannot be caught. Integrating [9] and [37], the exponential barrier function comes into effect if the nearest surrounding obstacle would collide with the robot, where the collision judgment is generated by the method in [37], as elaborated below.
@@
The switching mechanism is elucidated more clearly below: In the area $\mathcal{E}_{p}$, $\mu_{o}\neq 0$ in (23) is set to keep a safe distance from the obstacle, while in the region $\mathcal{E}_{c}$, we let $\mu_{o}=0$ to track the planned preliminary path. To satisfy the conditions C.1 and C.2 in the OMP Problem, the step cost function at the $k$-th time instant is re-designed as
@@ IV-D Online Receding-horizon Reinforcement Learning with a Game-based Exponential Barrier Function Design
Then, the optimal solution corresponding to the infinite-horizon optimization objective is defined by the following equation:
$\tilde{u}_{k}^{*}=\arg\min_{\tilde{u}_{k}}\sum_{\tau=k}^{\infty}\gamma^{\tau-k}L(\hat{\tilde{x}}_{\tau},\tilde{u}_{\tau}),$ (27)
@@
Note that the optimal solution and value function can be solved under the iterative learning framework[9]. However, the online convergence property is subject to the learning efficiency. Motivated by the concept of receding-horizon from MPC, as discussed in [11], we divide $V_{\infty}(\tilde{x}_{k})$ into multiple sub-problems within the prediction horizon $[k,k+N]$, where $N$ is the prediction horizon length. For any time instant $\tau$ in the prediction horizon, we have
[…]
@@
Eq. (28) is rewritten by
$V^{i+1}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\tilde{u}_{\tau})+\gamma V^{i}(\hat{\tilde{x}}_{\tau+1})\right],\tau\in[k,k+N-1],$ (29)
where the superscript of $V$ denotes the iteration number.
In alignment with the value function (25), one can derive its finite-horizon form by
$\lambda(\hat{\tilde{x}}_{\tau})=\partial V(\hat{\tilde{x}}_{\tau})/\partial\hat{\tilde{x}}_{\tau}.$ (30)
@@
For the critic network, the optimal value function corresponding to the finite-horizon optimization objective is defined by $\partial V^{*}(\hat{\tilde{x}}_{\tau})/\partial\hat{\tilde{x}}_{\tau}$, i.e,
@@
With the optimal value function $\lambda^{*}(\hat{\tilde{x}}_{\tau})$, one can compute the optimal solution by setting ${\partial V^{*}(\hat{\tilde{x}}_{\tau})}/{\partial\tilde{u}(\hat{\tilde{x}}_{\tau})}=0$. Then, for the actor network, the optimal solution corresponding to the finite-horizon optimization objective is defined by the following equation:
$\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})=-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\lambda^{*}(\hat{\tilde{x}}_{\tau+1}).$ (32)
@@
Output: The final safe trajecotry.
@@ 4 Function LPC:
5       Initialize the coefficient $\mu_{0}$ with 00.

### v4

@@ III-A Composite Vector Field
where $\hat{(\cdot)}$ is the normalization notation (i.e., for a nonzero vector $\in\mathbb{R}^{n}, v^=v$$\hat{v}$$\lVert v\rVert$), $\sqcup_{\mathcal{Q}}(\xi)=\frac{f_{1}(\xi)}{f_{1}(\xi)+f_{2}(\xi)},\sqcap_{\mathcal{R}}(\xi)=\frac{f_{2}(\xi)}{f_{1}(\xi)+f_{2}(\xi)}$ are smooth bump functions, where $f_{1}(\xi)=0$ if $\varphi(\xi)\leq c$ and $f_{1}(\xi)=\exp\left(\frac{l_{1}}{c-\varphi(\xi)}\right)$ if $\varphi(\xi)>c$, $f_{2}({\xi})=\exp\left(\frac{l_{2}}{\varphi(\xi)}\right)$ if $\varphi(\xi)<0$ and $f_{2}({\xi})=0$ if $\varphi(\xi)\geq 0$, and $l_{1},l_{2}>0$ are coefficients for changing the decaying or increasing rate. Note that for simplicity, the subscripts $i$ of related symbols have been omitted above. These smooth bump functions blend parts of different vector fields and create a composite vector field for path following and collision avoidance; for more details, see [5]. To understand the composite vector field intuitively, we illustrate the composite vector field (1) in Fig. 3. For $\chi(\xi)$ in (1), it is equal to $\hat{\chi}_{\mathcal{P}}(\xi)$, $\sqcup_{\mathcal{Q}}\left(\xi\right)\hat{\chi}_{\mathcal{P}}\left(\xi\right)+\sqcap_{\mathcal{R}}\left(\xi\right)\hat{\chi}_{\mathcal{R}}\left(\xi\right)$, and $\hat{\chi}_{\mathcal{R}}(\xi)$ within the three regions $^{\rm{ex}}{\mathcal{R}}$, $^{\rm{ex}}{\mathcal{Q}}\cap{}^{\mathrm{in}}\mathcal{R}$, and $^{\mathrm{in}}\mathcal{Q}$, respectively.
@@
$\mathcal{C}_{\mathcal{P}}$ $=\left\{\xi\in\mathbb{R}^{2}:\chi_{\mathcal{P}}(\xi)=0\right\}=\left\{\xi\in\mathbb{R}^{2}:\nabla\phi(\xi)=0\right\},$
$\mathcal{C}_{\mathcal{R}_{i}}$ $=\left\{\xi\in\mathbb{R}^{2}:\chi_{\mathcal{R}_{i}}(\xi)=0\right\}=\left\{\xi\in\mathbb{R}^{2}:\nabla\varphi_{i}(\xi)=0\right\}.$
@@ III-B Offline Deep Koopman Operator-based System Modeling
$\varUpsilon_{\infty}(x_{k+1})=(\mathcal{K}\varUpsilon_{\infty})(x_{k},u(x_{k})),$ (3)
@@ III-D2 Optimal Motion Planning to Avoid Moving Obstacles
$\varUpsilon(x_{k+1})=\underbrace{A\varUpsilon(x_{k})+Bu(x_{k})}_{f_{\text{nom}}(\varUpsilon(x_{k}),u(x_{k}))}+B_{s}\underbrace{\left(g(\varUpsilon(x_{k}),u(x_{k}))+w_{k}\right)}_{y_{k}},$ (9)
@@
$\tilde{x}_{k+1}=A\tilde{x}_{k}+B\tilde{u}(x_{k})+B_{s}\underbrace{\left(g(\varUpsilon(x_{k}),u(x_{k}))+w_{k}\right)}_{y_{k}},$ (11)
where $\tilde{x}_{k}=\varUpsilon(x_{k})-\varUpsilon(x_{r,k})$ is the error state, $x_{r,k}$ is the reference state, and $\tilde{u}(x_{k})=u(x_{k})-u(x_{r,k})$ is the control input.
@@
$V_{\infty}(\tilde{x}_{k})=\sum_{\tau=k}^{\infty}\gamma^{\tau-k}L(\tilde{x}_{\tau},\tilde{u}(\tilde{x}_{\tau})),$ (12)
where $0<\gamma\leq 1$, $L(\tilde{x}_{\tau},\tilde{u}_{\tau})=\tilde{x}_{\tau}^{\top}Q\tilde{x}_{\tau}+\tilde{u}^{\top}(\tilde{x}_{\tau})R\tilde{u}(\tilde{x}_{\tau})$ is the cost function, $Q\succeq 0\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$ is positive semi-definite, and $R\succ 0\in\mathbb{R}^{n_{u}\times n_{u}}$ is positive definite. The second subproblem is how the robot can avoid moving obstacles. Combining the two subproblems, we formulate the optimal OMP as below:
@@ IV-A1 Composite Vector Field with Kinodynamic Constraints
$\begin{aligned} \chi_{c}(\xi)=&\left(\prod_{i\in\mathcal{I}\backslash\mathcal{I}^{\prime}}\sqcup_{\mathcal{Q}_{i}}(\xi)\prod_{i\in\mathcal{I}^{\prime}}s_{i}(\xi)\right)\hat{\chi}_{\mathcal{P}}(\xi)\\ &+\sum_{i\in\mathcal{I}\backslash\mathcal{I}^{\prime}}\left(\sqcap_{\mathcal{R}_{i}}(\xi)\hat{\chi}_{\mathcal{R}_{i}}(\xi)\right)+\sum_{i\in\mathcal{I}^{\prime}}\left(1-{s_{i}}(\xi)\right)\hat{\chi}_{\mathcal{R}_{i}}(\xi),\end{aligned}$
@@
where the adjustable coefficient $k_{c,i}>0$ is used to change the convergence rate to $\mathcal{Q}_{\rm{vrtl}}$, and $\exp(\cdot)$ denotes the exponential function. Thus, in this design, $0<{s_{i}}(\xi)<1$ if $\xi\in\mathcal{M}_{b}$; ${s_{i}}(\xi)=1$, otherwise.
@@ IV-A3 Discrete-time Kinodynamic Guiding Vector Field
[…]
@@ IV-B Online Compensation to Update the Offline Trained Deep Koopman Model
${d}(z_{k})=B_{d}^{\dagger}(\varUpsilon({x}_{k+1})-f_{\text{nom}}\left(\varUpsilon({x}_{k}),{u}(x_{k})\right)),$ (18)
@@
$\varUpsilon({x}_{k+1})=f_{\text{nom}}(\varUpsilon({x}_{k}),{u}(x_{k}))+{d}(z_{k})+\epsilon,$ (19)
where $\epsilon\in\mathbb{R}^{n_{\mathcal{K}}}$ is the estimation error. The reference trajectory can be expressed as $\varUpsilon(x_{r,k+1})=f_{\text{nom}}(\varUpsilon({x}_{r,k}),{u}(x_{r,k}))$.
We can obtain the Jacobian matrix of (19) at a reference state $(\varUpsilon({x}_{r,k}),{u}(x_{r,k}))$, i.e.,
@@
$\hat{\tilde{x}}_{k+1}=f_{\text{nom}}(\tilde{x}_{k},\tilde{u}_{k}(\tilde{x}_{k}))+d({z_{k}}).$ (22)
@@ IV-C Exponential Barrier Function Incorporated Cost Function to Avoid Moving Obstacles
Due to the adoption of rule-based switching between tracking control and obstacle-avoiding modes in [9], without considering obstacle velocities, it does not guarantee the safety of mobile robots under extreme conditions such as the hazardous driving behavior of opposing vehicles. Our idea stems from the work[37] in completing suicidal game tasks, where the “evader” aspires to avoid being caught by the “pursuer” by computing safe actions. Similarly, when moving obstacles obstruct the road ahead, our strategy is to keep a safe distance from obstacles and track the desired trajectory if the ego robot cannot be caught. Integrating [9] and [37], the exponential barrier function comes into effect if the nearest surrounding obstacle would collide with the robot, where the judgment is generated by the method in [37], as elaborated below.
@@
The switching mechanism is elucidated more clearly below: In the area $\mathcal{E}_{p}$, $\mu_{o}\neq 0$ in (23) is set to keep a safe distance from the obstacle, while in the region $\mathcal{E}_{c}$, we let $\mu_{o}=0$ to track the planned preliminary path. To satisfy the conditions C.1 and C.2 in the OMP Problem, the step cost function at the $k$-th stage is re-designed as
@@ IV-D Online Receding-horizon Reinforcement Learning with a Game-based Exponential Barrier Function Design
Then, the optimal solution is defined by the following equation:
$\tilde{u}^{*}(\tilde{x}_{k})=\arg\min_{\tilde{u}(\tilde{x}_{k})}\sum_{\tau=k}^{\infty}\gamma^{\tau-k}L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau})),$ (27)
@@
Note that the optimal solution and value function can be solved under the iterative learning framework[9]. However, the online convergence property is subject to learning efficiency. Motivated by the concept of receding-horizon from MPC, as discussed in [11], we divide $V_{\infty}(\tilde{x}_{k})$ into multiple sub-problems within the prediction horizon $[k,k+N]$, where $N$ is the prediction horizon length. The value function in the prediction horizon is expressed by
$\left\{\begin{aligned} &V(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[F(\hat{\tilde{x}}_{k+N})+\sum_{j=\tau}^{k+N-1}L(\hat{\tilde{x}}_{j},\tilde{u}(\hat{\tilde{x}}_{j}))\right],\\ &V(\hat{\tilde{x}}_{k+N})=\mathbb{E}\left[F(\hat{\tilde{x}}_{k+N})\right],\end{aligned}\right$. (28)
@@
The value function $V(\hat{\tilde{x}}_{\tau})$ in Eq. (28) is written by
$V(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau}))+\gamma V(\hat{\tilde{x}}_{\tau+1})\right].$ (29)
According to the Bellman’s optimality principle, the optimal policy $\tilde{u}^{*}$ minimizes $V(\hat{\tilde{x}}_{\tau})$ in the prediction horizons; i.e.,
$\left\{\begin{aligned} &V^{*}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\tilde{u}^{*}(\hat{\tilde{x}}_{\tau}))+\gamma V^{*}(\hat{\tilde{x}}_{\tau+1})\right],\tau\in[k+N-1]\\ &V^{*}(\hat{\tilde{x}}_{k+N})=\mathbb{E}\left[F(\hat{\tilde{x}}_{k+N})\right].\end{aligned}\right$.
@@
For the critic network, the optimal value function in the finite horizon is re-defined by $\partial V^{*}(\hat{\tilde{x}}_{\tau})/\partial\hat{\tilde{x}}_{\tau}=0$, i.e,
@@
With the optimal value function $\lambda^{*}(\hat{\tilde{x}}_{\tau})$, one can compute the optimal solution by setting ${\partial V^{*}(\hat{\tilde{x}}_{\tau})}/{\partial\tilde{u}(\hat{\tilde{x}}_{\tau})}=0$. Then, for the actor network, the optimal solution in the finite horizon is defined by
$\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})=-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\lambda^{*}(\hat{\tilde{x}}_{\tau+1}),\tau\in[k,k+N-1].$ (32)
Remark 5.
The optimal solution $\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})$ and the optimal value function $\lambda^{*}(\hat{\tilde{x}}_{\tau})$ are difficult to be solved analytically. This is because $\lambda^{*}(\hat{\tilde{x}}_{\tau+1})$ is not available; see Eqs. (31) and (32). In the following, the optimal policies of the actor and critic networks are learned iteratively.
@@
Output: The final safe trajectory.
@@ 4 Function LPC:
5  Initialize the coefficient $\mu_{0}$ with 0.